# MCAQ-YOLO: Morphological Complexity-Aware Quantization for Efficient Object Detection with Curriculum Learning

### v1

@@ Abstract
Most neural network quantization methods apply uniform bit precision across spatial regions, ignoring the heterogeneous structural and textural complexity of visual data. This paper introduces MCAQ-YOLO, a morphological complexity-aware quantization framework for object detection. The framework employs five morphological metrics—fractal dimension, texture entropy, gradient variance, edge density, and contour complexity—to characterize local visual morphology and guide spatially adaptive bit allocation. By correlating these metrics with quantization sensitivity, MCAQ-YOLO dynamically adjusts bit precision according to spatial complexity. In addition, a curriculum-based quantization-aware training scheme progressively increases quantization difficulty to stabilize optimization and accelerate convergence. Experimental results demonstrate a strong correlation between morphological complexity and quantization sensitivity and show that MCAQ-YOLO achieves superior detection accuracy and convergence efficiency compared with uniform quantization. On a safety equipment dataset, MCAQ-YOLO attains 85.6% mAP@0.5 with an average of 4.2 bits and a 7.6 compression ratio, yielding 3.5 percentage points higher mAP than uniform 4-bit quantization while introducing only 1.8 ms of additional runtime overhead per image. Cross-dataset validation on COCO and Pascal VOC further confirms consistent performance gains, indicating that morphology-driven spatial quantization can enhance efficiency and robustness for computationally constrained, safety-critical visual recognition tasks.
@@
The deployment of deep neural networks for object detection in resource-constrained environments – ranging from edge devices to mobile platforms – necessitates aggressive model compression while maintaining high detection accuracy. Quantization, which reduces the numerical precision of weights and activations, has emerged as one of the most effective compression strategies, offering up to a theoretical speedup of for INT4 operations compared to FP32 on modern hardware [1, 2].
However, existing quantization approaches predominantly rely on layer-wise or channel-wise bit allocation schemes [3, 4], imposing uniform precision across spatial dimensions. Such uniformity overlooks a fundamental characteristic of visual data—the highly heterogeneous distribution of structural and textural complexity across spatial regions. In typical object detection scenarios, human instances with articulated poses, diverse clothing textures, and partial occlusions demand substantially higher representational precision than homogeneous backgrounds or geometrically simple objects. Empirical analysis confirms this trend: under mixed-precision quantization (3–6 bits), morphologically complex classes such as Person exhibit substantially larger performance degradation (15–18% mAP) when forced to lower bit-widths, whereas rigid objects such as helmets experience only modest degradation (4–6%, Mann–Whitney U test, , samples).
@@ I-A Motivation and Research Questions
This observation leads to the following central hypothesis: spatial regions exhibiting higher morphological complexity—such as irregular object boundaries, rich textures, and pronounced structural variations—require higher bit precision during quantization, whereas visually simple regions can be represented with more aggressive compression without incurring comparable losses in detection accuracy. This hypothesis is empirically motivated rather than derived from formal information-theoretic principles and therefore warrants systematic validation.
To examine this hypothesis, we investigate the following research questions:
@@ I-B Technical Contributions
The main contributions of this work are summarized as follows:
@@
Multi-scale morphological complexity framework: We develop a morphological complexity analyzer that combines five complementary descriptors with an efficient multi-scale implementation. The proposed hierarchical approximation reduces the computational cost of fractal-dimension estimation from to through multi-resolution box counting with logarithmic scale sampling, while empirically maintaining a small relative error (typically ).
@@
Curriculum learning for quantization-aware training: We incorporate a curriculum learning scheme into quantization-aware training by progressively increasing the complexity of training samples. On the main benchmark, this yields 2.5× faster convergence to reach comparable performance levels (20k versus 50k iterations) and roughly 60% lower gradient variance, indicating more stable optimization.
@@
Learnable complexity-to-bit mapping: We design a differentiable mapping network that learns the non-linear relationship between morphological features and spatial bit allocation and empirically reveals phase-transition-like behaviors in the complexity–precision trade-off.
@@
Statistically grounded evaluation: We conduct extensive experiments using rigorous statistical methodology, including Benjamini–Hochberg false-discovery-rate control for multiple comparisons, bootstrap confidence intervals (10,000 resamples), and post-hoc power analysis confirming approximately 95% power for the observed effects.
- 5.
@@ I-C Scope, Assumptions, and Limitations
Scope: This work focuses on spatial quantization within single-stage object detectors, specifically the YOLO family. We target inference optimization rather than training efficiency.
Assumptions:
Hardware supports layer-wise mixed precision (validated on NVIDIA TensorRT 8.6). Tile-wise precision is approximated through grouped convolutions with different bit-widths per group.
@@
Limitations acknowledged upfront:
The current unoptimized implementation adds 16.4 ms (CPU) or 4.5 ms (GPU) overhead, which reduces to 1.8 ms with optimizations enabled.
Theoretical justification for input-based bit allocation remains an open problem.
Performance gains vary by dataset characteristics (approximately 3.5% on datasets with high morphological variability and around 2–3% on datasets with more uniform complexity).
@@ II-A Neural Network Quantization
Quantization maps continuous values to discrete representations and has been widely studied for both post-training quantization (PTQ) and quantization-aware training (QAT). PTQ methods, such as BRECQ [8] and AdaRound [9], reconstruct layer or block outputs of pretrained models and optimize rounding or scaling decisions without full retraining. QAT methods, including LSQ [10] and PACT [11], introduce differentiable approximations of quantization operators during training, allowing scales and clipping ranges to be learned jointly with network weights.
Mixed-precision quantization further assigns different bit-widths to different layers or channels. Representative approaches include HAQ [13], which employs reinforcement learning with hardware feedback, and the HAWQ series [3, 14, 4], which leverages Hessian information to quantify layer sensitivity. More recently, information-entropy-based schemes have been proposed to allocate bits according to activation entropy at the layer level [6]. However, these methods predominantly operate at layer or channel granularity and do not explicitly account for spatial variations within feature maps, which is the focus of MCAQ-YOLO.
@@ II-B Mathematical Morphology and Complexity Metrics
Mathematical morphology provides tools to quantify geometric and textural characteristics of images. The fractal dimension [15] is frequently used as a measure of geometric complexity. In practice, the box-counting dimension is widely adopted:
@@
In MCAQ-YOLO, we combine fractal dimension, entropy-based texture descriptors, gradient statistics, edge density, and contour-based shape descriptors into a unified morphological complexity score that directly drives spatial bit allocation.
@@ II-C Curriculum Learning
Curriculum learning [19] trains models on progressively harder examples by controlling the distribution of samples presented during optimization. This strategy has been applied to neural architecture search [20], self-paced learning for detection [21], and video understanding [22]. To our knowledge, MCAQ-YOLO is the first to apply curriculum learning to quantization-aware training, where the “difficulty” of samples is defined in terms of morphological complexity and the induced quantization sensitivity.
@@ III Theoretical Framework and Analysis
III-A Empirical Hypothesis: Morphological Complexity and Quantization Sensitivity
We adopt an observation-driven framework that links morphological complexity to quantization sensitivity.
@@ III-A Empirical Hypothesis: Morphological Complexity and Quantization Sensitivity
We validate these hypotheses through correlation analysis between morphological complexity and quantization-induced performance degradation and through ablation studies on bit allocation policies.
III-B Morphological Complexity Formulation
@@
To better capture interactions between different aspects of complexity, we use an enhanced score:
@@
III-C Quantization Error and Detection Performance
For uniform -bit quantization, the spatial distribution of quantization error depends on local statistics of feature values. We model the expected squared error at position as
@@ III-C Quantization Error and Detection Performance
III-D Computational Complexity of Morphological Analysis
@@ IV MCAQ-YOLO Architecture and Implementation
IV-A System Architecture
MCAQ-YOLO augments a baseline YOLOv8 detector with three key modules: a hierarchical morphological complexity analyzer that operates on intermediate feature maps, a learnable mapping network that converts complexity scores into spatially varying bit allocations, and a hardware-aware quantization module that realizes tile-wise mixed precision while preserving smooth transitions between regions. These components are trained jointly within a multi-objective framework that balances detection accuracy, bit budget, and spatial smoothness.
@@ IV-B Hierarchical Morphological Complexity Analyzer
The analyzer computes morphological complexity over tiles of intermediate feature maps using a multi-scale, cache-aware algorithm summarized in Algorithm 1.
@@ IV-C Curriculum-Trained Complexity-to-Bit Mapping Network
The mapping network learns a function that converts local complexity scores into effective bit-widths:
To encourage monotonicity, we constrain weights to be non-negative by taking absolute values during training:
Curriculum learning is applied by gradually increasing the maximum allowed complexity of training samples and by annealing the “temperature” of bit allocation. The overall procedure is summarized in Algorithm 3.
@@ IV-D Hardware-Aware Smooth Transition Quantization
For hardware efficiency, we implement tile-wise mixed-precision quantization with non-overlapping tiles. Let denote the index of the tile that contains spatial position . The quantized feature at is given by
where is a learned soft mask that smoothly varies over space and denotes quantization with bit-width for tile . Tiles are non-overlapping, and each spatial position belongs to exactly one tile; smooth transitions between regions are obtained by spatially smoothing rather than by summing contributions from multiple tiles. The masks are produced by a softmax-based module followed by spatial smoothing so that changes gradually across neighboring tiles, which avoids visible artifacts at tile boundaries while remaining compatible with a single-tile assignment per spatial location in the fused CUDA kernel.

### v2

@@ Abstract
Most neural network quantization methods apply uniform bit precision across spatial regions, disregarding the heterogeneous complexity inherent in visual data. This paper introduces MCAQ-YOLO, a practical framework for tile-wise spatial mixed-precision quantization in real-time object detectors. Morphological complexity—quantified through five complementary metrics (fractal dimension, texture entropy, gradient variance, edge density, and contour complexity)—is proposed as a signal-centric predictor of spatial quantization sensitivity. A calibration-time analysis design enables spatial bit allocation with only 0.3 ms inference overhead, achieving 151 FPS throughput. Additionally, a curriculum-based training scheme that progressively increases quantization difficulty is introduced to stabilize optimization and accelerate convergence. On a construction safety equipment dataset exhibiting high morphological variability, MCAQ-YOLO achieves 85.6% mAP@0.5 with an average bit-width of 4.2 bits and a 7.6 compression ratio, outperforming uniform 4-bit quantization by 3.5 percentage points. Cross-dataset evaluation on COCO 2017 (+2.9%) and Pascal VOC 2012 (+2.3%) demonstrates consistent improvements, with performance gains correlating with within-image complexity variation.
@@
Deploying deep neural networks for object detection in resource-constrained environments—ranging from edge devices to mobile platforms—necessitates aggressive model compression while maintaining high detection accuracy. Quantization, which reduces the numerical precision of weights and activations, has emerged as one of the most effective compression strategies, offering up to 16 theoretical speedup for INT4 operations compared to FP32 on modern hardware [1, 2].
However, existing quantization approaches predominantly employ layer-wise or channel-wise bit allocation schemes [3, 4], imposing uniform precision across spatial dimensions. Such uniformity overlooks a fundamental characteristic of visual data: the highly heterogeneous distribution of structural and textural complexity across spatial regions. In typical object detection scenarios, human instances with articulated poses, diverse clothing textures, and partial occlusions demand substantially higher representational precision than homogeneous backgrounds or geometrically simple objects. Empirical analysis confirms this observation: under mixed-precision quantization (3–6 bits), morphologically complex classes such as Person exhibit substantially larger performance degradation (15–18% mAP) when constrained to lower bit-widths, whereas rigid objects such as helmets experience only modest degradation (4–6%; Mann–Whitney U test, , samples).
@@ I-A Motivation and Research Questions
This observation motivates our central hypothesis: spatial regions exhibiting higher morphological complexity—characterized by irregular object boundaries, rich textures, and pronounced structural variations—require higher bit precision during quantization, whereas visually simple regions can be represented with more aggressive compression without incurring comparable losses in detection accuracy.
A new perspective on quantization sensitivity. While prior work has characterized layer-wise or channel-wise sensitivity using Hessian traces or activation entropy, this work proposes that morphological complexity—a property of the input signal itself—serves as a direct and interpretable predictor of spatial quantization sensitivity. This perspective represents a shift from network-centric metrics to signal-centric metrics, enabling spatially adaptive bit allocation at a granularity previously unexplored in practical object detection systems.
To investigate this hypothesis, we address the following research questions:
@@ I-B Technical Contributions
The principal contributions of this work are summarized as follows:
@@
Morphological complexity as a quantization sensitivity predictor: This work formalizes the relationship between visual morphology and quantization sensitivity, proposing that signal-level complexity metrics—fractal dimension, texture entropy, gradient variance, edge density, and contour complexity—provide a principled basis for spatial bit allocation. Unlike layer-wise Hessian methods [3] or channel-wise entropy approaches [6], the proposed formulation operates at spatial tile granularity, enabling fine-grained precision control within feature maps.
@@
Practical tile-wise spatial mixed-precision quantization: While spatial mixed-precision has been theoretically appealing, deploying it in real-time detectors has remained challenging due to inference overhead. This gap is bridged through a calibration-time analysis with inference-time lookup design: morphological analysis runs once during calibration to precompute bit-maps, enabling tile-wise quantization at inference with only 0.3 ms additional latency (151 FPS on RTX PRO 6000). To the best of our knowledge, this represents the first practical realization of spatial mixed-precision quantization for real-time object detection.
@@
Curriculum learning for quantization-aware training: A curriculum learning scheme is incorporated into quantization-aware training by progressively increasing the complexity of training samples. On the main benchmark, this approach yields faster convergence to comparable performance levels (20k versus 50k iterations) and approximately 60% lower gradient variance, indicating more stable optimization.
@@
Statistically grounded evaluation: Extensive experiments are conducted using rigorous statistical methodology, including Benjamini–Hochberg false-discovery-rate control for multiple comparisons, bootstrap confidence intervals (10,000 resamples), and post-hoc power analysis confirming approximately 95% power for the observed effects.
@@ I-C Scope, Assumptions, and Limitations
Scope. This work focuses on spatial quantization within single-stage object detectors, specifically the YOLO family. We target inference optimization rather than training efficiency.
Assumptions.
Hardware supports layer-wise mixed precision (validated on NVIDIA TensorRT 8.6). Tile-wise precision is realized through a custom CUDA kernel (see Sec. IV-D).
@@
Limitations.
Empirical foundation: The complexity–sensitivity relationship is empirically established rather than theoretically grounded; formal justification remains an open problem.
Hardware deployment: The current unoptimized implementation adds 16.4 ms (CPU) or 4.5 ms (GPU) overhead, which reduces to 1.8 ms with all optimizations enabled. Production deployment may require further kernel optimization.
Dataset dependency: Performance gains vary with dataset characteristics—approximately 3.5% on datasets with high morphological variability and 2–3% on datasets with more uniform complexity.
Architecture scope: Validation is limited to CNN-based YOLO detectors; extensions to Transformer-based architectures (e.g., DETR, RT-DETR) require additional investigation.
@@ II-A Neural Network Quantization
Quantization maps continuous values to discrete representations and has been widely studied for both post-training quantization (PTQ) and quantization-aware training (QAT). PTQ methods, such as BRECQ [9] and AdaRound [10], reconstruct layer or block outputs of pretrained models and optimize rounding or scaling decisions without full retraining. QAT methods, including LSQ [11] and PACT [12], introduce differentiable approximations of quantization operators during training, enabling scales and clipping ranges to be learned jointly with network weights.
Mixed-precision quantization assigns different bit-widths to different layers or channels. Representative approaches include HAQ [14], which employs reinforcement learning with hardware feedback, and the HAWQ series [3, 15, 4], which leverages Hessian information to quantify layer sensitivity. More recently, information-entropy-based schemes have been proposed to allocate bits according to activation entropy at the layer level [6].
Recent advances in large language model compression have introduced efficient quantization techniques such as GPTQ [7], which achieves accurate 3–4 bit quantization through one-shot weight quantization with approximate second-order information, and QLoRA [8], which enables fine-tuning of quantized models through low-rank adapters. While these methods primarily target language models, they demonstrate the viability of aggressive quantization with minimal accuracy loss.
However, these methods predominantly operate at layer or channel granularity and do not explicitly account for spatial variations within feature maps, which is the focus of MCAQ-YOLO.
@@ II-B Mathematical Morphology and Complexity Metrics
Mathematical morphology provides tools to quantify geometric and textural characteristics of images. The fractal dimension [16] is frequently used as a measure of geometric complexity via box-counting. In practice, the box-counting dimension is widely adopted:
@@
In MCAQ-YOLO, fractal dimension, entropy-based texture descriptors, gradient statistics, edge density, and contour-based shape descriptors are combined into a unified morphological complexity score that directly drives spatial bit allocation.
@@ II-C Curriculum Learning
Curriculum learning [20] trains models on progressively harder examples by controlling the distribution of samples presented during optimization. This strategy has been applied to neural architecture search [21], self-paced learning [22], and video understanding [23]. To the best of our knowledge, MCAQ-YOLO is the first work to apply curriculum learning to quantization-aware training, where sample “difficulty” is defined in terms of morphological complexity and the induced quantization sensitivity.
@@ II-D Object Detection Architectures
II-E Distinction from Entropy-Based Mixed Precision
While entropy-driven quantization methods [6] allocate bits based on activation entropy at the layer or channel level, MCAQ-YOLO operates at a fundamentally different granularity: spatial tiles within feature maps. This distinction is critical for object detection, where a single feature map may contain both simple backgrounds and complex object boundaries requiring different precision levels. Furthermore, the proposed morphological complexity score incorporates geometric structure (fractal dimension, contour complexity) that pure entropy measures cannot capture. Activation entropy alone exhibits weaker correlation with quantization sensitivity (, tile-level) compared to the unified complexity score (), particularly for boundary-rich regions. The learnable complexity-to-bit mapping and curriculum-based training further distinguish this approach by adapting the complexity–precision relationship to specific detection tasks rather than relying on fixed entropy thresholds.
@@ III Theoretical Framework and Analysis
III-A Rate–Distortion Perspective
To provide additional mathematical grounding for why morphologically complex regions should receive more bits, we draw on rate–distortion theory. Consider a local feature tile (or spatial region) represented by a random variable and its quantized reconstruction . The minimum number of bits required to achieve an expected distortion is characterized by the rate–distortion function:
where denotes mutual information and is typically mean squared error (MSE). Under a Gaussian approximation for local signals, this simplifies to:
implying that for a fixed distortion budget , higher local variance demands a higher coding rate.
Connection to morphological complexity. Morphologically complex regions—edges, textures, and irregular contours—exhibit properties that increase their effective rate demand: (i) higher local variance due to rapid intensity transitions, (ii) broader frequency content requiring finer quantization to preserve high-frequency components, and (iii) greater sensitivity to quantization noise in downstream detection, as boundary perturbations directly affect localization accuracy. The proposed morphological descriptors (fractal dimension, gradient variance, edge density) serve as computationally efficient proxies for these information-theoretic quantities, obviating the need for explicit entropy estimation at each spatial location.
For a uniform scalar quantizer with step size , the quantization noise variance is . Since for a fixed dynamic range, reducing distortion necessitates increasing bit-width . MCAQ-YOLO instantiates this principle by learning a mapping from morphology-derived complexity to spatially adaptive bit allocation under a global bit budget.
III-B Empirical Hypothesis: Morphological Complexity and Quantization Sensitivity
An observation-driven framework is adopted that links morphological complexity to quantization sensitivity.
@@ III-B Empirical Hypothesis: Morphological Complexity and Quantization Sensitivity
As illustrated in Fig. 1, regions with higher morphological complexity tend to produce activation distributions with sharper transitions and higher-frequency components. Such distributions are known to be more sensitive to low-bit quantization, leading to disproportionately large quantization errors under uniform bit-width allocation. This observation motivates allocating additional bits to high-complexity regions, while aggressively reducing precision in low-complexity regions with minimal performance degradation.
These hypotheses are validated through correlation analysis between morphological complexity and quantization-induced performance degradation and through ablation studies on bit allocation policies.
III-C Morphological Complexity Formulation
@@
To better capture interactions between different aspects of complexity, an enhanced score is employed:
@@
III-D Quantization Error and Detection Performance
For uniform -bit quantization, the spatial distribution of quantization error depends on local statistics of feature values. The expected squared error at position is modeled as
@@ III-D Quantization Error and Detection Performance
III-E Computational Complexity of Morphological Analysis
@@ IV MCAQ-YOLO Architecture and Implementation
IV-A System Overview
MCAQ-YOLO augments a baseline YOLOv8 detector with three key modules: (1) a hierarchical morphological complexity analyzer that operates on intermediate feature maps, (2) a learnable mapping network that converts complexity scores into spatially varying bit allocations, and (3) a hardware-aware quantization module that realizes tile-wise mixed precision while preserving smooth transitions between regions. These components are trained jointly within a multi-objective framework that balances detection accuracy, bit budget, and spatial smoothness. The overall architecture is illustrated in Fig. 2.
@@ IV-B Hierarchical Morphological Complexity Analyzer
The analyzer computes five complementary descriptors on intermediate feature maps and aggregates them into a unified complexity score as defined in Eq. (8). Each descriptor is computed efficiently using optimized implementations:
Fractal dimension is estimated using a fast multi-resolution box-counting algorithm that samples dyadic scales, achieving complexity through efficient max-pooling operations.
Texture entropy is computed from LBP histograms over local neighborhoods, capturing fine-grained textural patterns that correlate with quantization sensitivity.
Gradient variance measures local contrast and edge strength using Sobel operators, providing a direct indicator of high-frequency content.
Edge density quantifies the proportion of edge pixels within each tile using Canny edge detection with adaptive thresholds.
Contour complexity encodes boundary irregularity through shape descriptors including perimeter-to-area ratio and Fourier descriptor statistics.
To reduce computational overhead, a caching mechanism is employed that stores complexity scores for tiles with similar feature statistics (determined by locality-sensitive hashing). Training note: The morphological descriptors are computed as deterministic side-information; gradients are propagated through the mapping network and quantization operator (via straight-through estimator), not through the morphology computation itself.
The complete algorithm with optimization details is provided in Algorithm 1.
AdaptiveTileSize. The tile size is selected based on expected object density. For static deployment, density is estimated from calibration set statistics. For video streams, the previous frame’s detection count is used (one-frame delay). For single novel images, a default grid is applied:
Caching mechanism. The locality-sensitive hash function computes a compact signature from downsampled tile statistics (mean, variance, histogram bins). For video/streaming applications where consecutive frames share similar backgrounds, this enables cache lookup with approximately 85% hit rate. For static image datasets (as in Table II), caching is not applicable; instead, bit-maps are precomputed during calibration and stored for inference.
@@ IV-C Curriculum-Trained Complexity-to-Bit Mapping Network
The mapping network learns a function that converts local complexity scores into effective bit-widths. The network architecture incorporates polynomial feature expansion to capture nonlinear relationships:
To encourage monotonicity (higher complexity higher bits), we constrain weights to be non-negative by taking absolute values during training:
Curriculum learning is applied by gradually increasing the maximum allowed complexity of training samples and by annealing the “temperature” of bit allocation. The complexity threshold increases linearly from to during the warmup phase, while the temperature decays exponentially from 10 to 1. The three-stage curriculum schedule is illustrated in Fig. 3:
Stage 1 (Epochs 0–20): Warm-up with low-complexity samples only () and high precision (FP16).
Stage 2 (Epochs 20–50): Transition with mixed-complexity samples and dynamic bit allocation; temperature annealing begins.
Stage 3 (Epochs 50–300): Full MCAQ with all samples and aggressive quantization; fine-tuning for details.
The curriculum learning algorithm is detailed in Algorithm 3.
@@ IV-D Hardware-Aware Smooth Transition Quantization
For hardware efficiency, tile-wise mixed-precision quantization is implemented using a custom CUDA kernel that performs per-element quantization with spatially varying bit-widths. The spatial quantization is applied at the output of the backbone (C3/C4/C5 feature maps) before the FPN neck, where feature resolution is sufficient for meaningful tile-level complexity variation while computational overhead remains manageable.
The feature map is partitioned into non-overlapping tiles, and each tile is assigned a bit-width determined by the complexity-to-bit mapping network. The quantized feature at position is given by:
where is a learned soft mask that varies smoothly over space and denotes quantization with bit-width for tile . Tiles are non-overlapping, and each spatial position belongs to exactly one tile; smooth transitions between regions are obtained by spatially smoothing rather than by summing contributions from multiple tiles. The masks are produced by a softmax-based module followed by spatial smoothing so that changes gradually across neighboring tiles, which avoids visible artifacts at tile boundaries while remaining compatible with a single-tile assignment per spatial location in the fused CUDA kernel.
The tile-wise quantization processes each spatial position independently: it determines the tile index from spatial coordinates, retrieves the corresponding bit-width from the precomputed bit map, and applies symmetric uniform quantization with dynamically computed scale and zero-point. This element-wise implementation avoids the overhead of grouped convolutions and enables true spatial mixed precision at tile granularity, with each tile operating at its assigned bit-width (2–8 bits). The CUDA kernel implementation is provided in Appendix A.
Calibration and tile configuration. Per-channel min/max statistics are collected from 1,000 calibration images using exponential moving average (EMA) with momentum 0.99, then frozen to compute scale and zero-point per channel. The tile size defaults to (i.e., an grid for inputs), with finer grids used adaptively for high object density scenarios.